Assume you buy a bond with the following features Bond maturity = 4 Coupon Rate = 5.00% Face Value = $1,000 Annual Coupons When you buy the bond the market interest rate = 6.00% Immediately after you buy the bond the interest rate changes to 5.50% What is the "reinvestment" effect in year 3 ?
Answer #1

Bond Tenure = 4 years, Coupon Rate = 5 %, Face Value = $ 1000, Initial Market Interest Rate = 6 % and Changed Market Interest Rate = 5.5 %

Annual Coupon = 0.05 x 1000 = $ 50

Reinvested Coupon Value at the end of Year 3 at 6% = 50 x (1.06)^(3) + 50 x (1.06) + 50 = $ 159.18

Reinvested Coupon Value at the end of Year 3 at 5.5 % = 50 x (1.055)^(2) + 50 x (1.055) + 50 = $ 158.40125

Reinvestment Effect = 158.40125 - 159.18 = - $ 0.77875